OpenStudy (anonymous):

Simplify 13x - 4[x + (3 - x)]. a. 13x - 12 b. 9x - 1 c. 8x - 12

OpenStudy (anonymous):

the answer is a i think

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

2abc + d = 3 ?

OpenStudy (anonymous):

solve for c

OpenStudy (anonymous):

its eithor c = (3-d)/(2ab) or 3/(2ab) -d sorry thats the best i can do on this one

OpenStudy (anonymous):

it was the first one. thanks

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

What is the range of the function y = -|x|?

OpenStudy (anonymous):

\[x \le 0... or... y \le0... or... all real numbers\]

OpenStudy (anonymous):

i think its all real numbers..

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (-1, -1) and is parallel to the line whose equation is x + y = 3?

OpenStudy (anonymous):

what are you choices? for answers?

OpenStudy (anonymous):

@Jadeishere

OpenStudy (anonymous):

y = - 1/2x + 1 y = 1/2x - 1 y = 2x + 1

OpenStudy (anonymous):

^^^^choices for answers

OpenStudy (anonymous):

sorry i really dont know this one if i would take a guess it would be y = -1/2x +1 but thats just a guess

OpenStudy (anonymous):

thnx anyways

OpenStudy (anonymous):

no problem any other questions?

OpenStudy (anonymous):

soon

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 2 y + 4 = 3x 3x + 3x - 4 = 2 3x + 3 - 4x = 2 3x + 3x + 4 = 2

OpenStudy (anonymous):

ill tey to answer that in a little bit i got to gof for like 30 minnetes sorry

OpenStudy (anonymous):

ugh ok.